OpenStudy (anonymous):

The half-time for the elimination of a drug is h, show that k=ln2/h. Where A=De^(kt).

OpenStudy (anonymous):

y-2x-6=0 and y/2-3=x solve the two variables

OpenStudy (anonymous):

A friend of mine did it like this D=De^kh D/2=De^kh 1/2=e^kh ln(1/2)=kh (ln(1/2))/h=k And somehow got k=(ln2)/ h

OpenStudy (anonymous):

Wait, k is supposed to be negative.

OpenStudy (anonymous):

please help me solve y-2x-6=and y/2-3=x

OpenStudy (anonymous):

wtf? Make your own question haha

OpenStudy (callisto):

I think it is A=De^(-kt) in the question?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

But how would you get k=ln2/h?

OpenStudy (callisto):

Now,that makes sense A=De^(-kt), where A = D/2, t=h A=De^(kt) D/2 =De^(-kh) 1/2 = e^(-kh) Take ln on both sides ln (1/2) = ln e^(-kh) -ln2 = (-kh) lne , note that lne =1 -ln2 = -kh k = (ln2)/h Does it make sense to you?

OpenStudy (anonymous):

The step where ln(1/2) goes to -ln2 is confusing me

OpenStudy (anonymous):

Oh wait ln(1/2) is equal to -ln2?

OpenStudy (callisto):